# On balanced Hermitian threefolds with parallel Bismut torsion

### v1

@@ 3. The Fano case
[…]
@@
It follows from [35, Theorem 1.1] that the BTP condition $\nabla^{b}T=0$ implies
@@
That is, $\tilde{\omega}$ is a Kähler-Einstein metric with positive Ricci curvature $2$, hence $M^{3}$ is a Fano threefold. Denote by $E$ and $F$ the $C^{\infty}$ complex vector bundles on $M$ whose fibers are $E_{x}={\mathbb{C}}\{e_{2}(x),e_{3}(x)\}$ and $F_{x}={\mathbb{C}}e_{1}(x)$, respectively. They are both globally defined since $E$ is the eigenspace of $B$ corresponding to the eigenvalue 00, and $F$ is the orthogonal complement of $E$ in the holomorphic tangent bundle $T_{\!M}=T^{1,0}M$. For any $i\in\{2,3\}$ and any $j$, we have
@@
(3.7) $0\rightarrow E\rightarrow T_{\!M}\rightarrow L\rightarrow 0$
@@
Recall that the index of a Fano manifold $X^{n}$ is the largest positive integer $r$ so that $K^{\!-\!1}_{\!X}=A^{\otimes r}$ for an ample line bundle $A$. It is necessarily less than or equal to $n+1$, where $r=n+1$ if and only if $X={\mathbb{P}}^{n}$ and $r=n$ if and only if $X={\mathbb{Q}}^{n}$, the smooth quadratic hypersurface in ${\mathbb{P}}^{n+1}$. Fano manifolds satisfying $r=n-1$ are called del Pezzo manifolds, which are classified by Fujita [17] as one of the following seven types according to their degree $d$, which is the self intersection number $A^{n}$:
@@ 7. $(7)$
For $n=3$, del Pezzo threefolds were classified by Iskovskikh [18] earlier, and in Table 12.2 of [19] we can find the third betti number $b_{3}$, hence the Euler number $c_{3}=4-b_{3}$ of del Pezzo threefolds of degree $1\leq d\leq 5$:
@@
If $M^{3}$ is a del Pezzo threefold of degree $d$, then $c_{1}^{3}=8d$, hence the equality (3.10) yields $c_{3}=12-d$. This rules out the possibility of $1\leq d\leq 5$ by (3.9). The case ${\mathbb{P}}^{3}\#\overline{{\mathbb{P}}^{3}}$ of degree $d=7$ has Euler number $c_{3}=6$, which is not equal to $12-7$. Similarly, for the case ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ of degree $d=6$, its Euler number $c_{3}=8$ is not equal to $12-6$, which indicates that neither can be $M^{3}$. Therefore only two possibilities are left, namely, $M^{3}$ is either the flag threefold ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$ or ${\mathbb{P}}^{3}$.
@@
To see this, let $e$ be a local special frame of $(M^{3},g)$, with dual coframe $\varphi$. Let $s$ be a local holomorphic frame of $L$. Since $\xi$ is a $L$-valued $1$-form, we have $\xi=\psi\otimes s$ where $\psi$ is a nowhere-zero local holomorphic $1$-form. The kernel of the map given by $\xi$ is $E$, so $\psi=f\varphi_{1}$ for some local smooth function $f$, which is nowhere zero. By the structure equation, it follows that $d\varphi_{1}=-(x+y)\varphi_{1}-\varphi_{2}\varphi_{3}$, so the holomorphicity of $\psi$ gives us
@@
(3.11) ¯$\partial\overline{\partial}\log(|f|^{2}\!\parallel\!\!s\!\!\parallel^{2})=0.$
@@
Assume that $M^{3}={\mathbb{P}}^{3}$. In this case $L={\mathcal{O}}(2)$. It follows from [26] for instance that $\Omega(2)$ admits a trivial line subbundle, which corresponds to a nowhere zero holomorphic section $\xi\in V:=H^{0}({\mathbb{P}}^{3},\Omega(2))$. $\xi$ determines a surjective bundle map $T_{\mathbb{P}^{3}}\rightarrow L$ so that the exact sequence (3.7) holds. Let $\tilde{\omega}$ be the (scaled) Fubini-Study metric of ${\mathbb{P}}^{3}$ with Ricci curvature $2$. It has constant holomorphic sectional curvature $1$. Let $[Z_{0}\!:\!Z_{1}\!:\!Z_{2}\!:\!Z_{3}]$ be the standard unitary homogeneous coordinate of ${\mathbb{P}}^{3}$. In the coordinate neighborhood $U_{0}=\{Z_{0}\neq 0\}$, let $z_{i}=\frac{Z_{i}}{Z_{0}}$, $1\leq i\leq 3$ and it follows that
@@
where $c>0$ is a constant. We want to derive at a contradiction. For each $1\leq i\leq 3$, write $\ell_{i}=\ell_{i}^{(0)}+\ell_{i}^{(1)}$ where $\ell_{i}^{(0)},\ \ell_{i}^{(1)}$ are respectively the degree 00 and degree $1$ part of $\ell_{i}$. By looking at the degree $4$ part of the left hand side of (3.12), we get
@@
Now we are left with the only possibility for a compact balanced BTP threefold $(M^{3},g)$ with $B$ rank equal to $1$: the flag threefold $X^{3}={\mathbb{P}}(T_{{\mathbb{P}}^{2}})$. Note that $X^{3}$ is the hypersurface in $N^{4}={\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ defined by $Z_{0}W_{0}+Z_{1}W_{1}+Z_{2}W_{2}=0$, where $Z$, $W$ are the standard unitary homogeneous coordinate of the two factors of $N$. For $i=1,2$, denote by $\pi_{i}:X^{3}\rightarrow{\mathbb{P}}^{2}$ the restriction on $X$ of the projection map from $N$ onto its $i$-th factor. The Picard group $\mbox{Pic}(X)\cong{\mathbb{Z}}^{\oplus 2}$ is generated by $L_{1}$ and $L_{2}$, where $L_{i}=\pi_{i}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)$, and the anti-canonical line bundle of $X$ is $-K_{\!X}=2L=2(L_{1}+L_{2})$. Here for convenience we used the additive notation for line bundles. The Kähler-Einstein metric $\tilde{\omega}$ on $X$ is the restriction of the product of Fubini-Study metric, and we have
@@
Let $X^{3}$ be the flag threefold ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$, $\Omega$ its holomorphic cotangent bundle, and $L$ be the line budle $\pi_{1}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)\otimes\pi_{2}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)$. Then we have
@@
$0\rightarrow T_{X|{\mathbb{P}}^{2}}\rightarrow T_{X}\rightarrow\pi_{1}^{\ast}T_{{\mathbb{P}}^{2}}\rightarrow 0$
@@
(3.15) $0\rightarrow{\mathcal{O}}_{X}\rightarrow\pi_{1}^{\ast}\Omega_{{\mathbb{P}}^{2}}\!\otimes\!L\rightarrow T_{X|{\mathbb{P}}^{2}}\rightarrow 0$
@@
where $Z$, $W$ are unitary homogeneous coordinate on the two factors of $N={\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$, viewed as column vectors. It is not hard to see that the norm $\parallel\!\!\sigma\!\!\parallel=\frac{1}{2}$ with respect to the Kähler-Einstein metric $\tilde{\omega}$ of $X$. Consider the Hermitian metric $g$ on $X$ with Kähler form $\omega=\tilde{\omega}-\sigma$, which is clearly a homogeneous Hermitian metric on $X$. We will call the Hermitian manifold $(X,g)$ the Wallach threefold from now on, to honor the influential work [31] in geometry.
We will verify in the next section that $(X,g)$ is indeed balanced and BTP. We will also show that its Chern connection has non-negative bisectional curvature and positive holomorphic sectional curvature, and all three Ricci tensors of the Chern connection are positive. The sectional curvature of the Levi-Civita connection of $g$ is non-negative, and the Levi-Civita connection has constant Ricci curvature $3$, thus $g$ lies in the boundary of the set of metrics with positive sectional curvature discovered by Wallach in [31].
Note that homogeneous metrics on $X$ with positive sectional curvature, which are all Hermitian as observed by Wallach in [31], form a moduli which depends on three real parameters. After scaling, these metrics form a peculiar planer region (see for example Figure 1 in [11]). It is not clear which metric in the set is the ‘best’ amongst its peers.
@@ 4. The Wallach threefold
Fix any point $p\in X=SU(3)/T^{2}$. Note that for any $A\in SU(3)$, the map $([Z],[W])\mapsto([AZ],[\overline{A}W])$ is an isometry on $(X,g)$. So without loss of generality, we may assume that $p=([1\!\!:\!0\!:\!0],[0\!:\!0\!:\!\!1])$. Consider the neighborhood $U_{02}=\{Z_{0}\neq 0\}\times\{W_{2}\neq 0\}$ in $N$, with local holomorphic coordinate $(z_{1},z_{2},w_{0},w_{1})$ where $z_{i}=\frac{Z_{i}}{Z_{0}}$, $i=1,2$, and $w_{j}=\frac{W_{j}}{W_{2}}$, $j=0,1$. Within $U_{02}$, the hypersurface $X$ is defined by
@@
[…]
@@
(4.5) $\tilde{g}_{i\overline{j},k}=-\frac{1}{\alpha^{2}}\big{(}\alpha_{k}\alpha_{i\overline{j}}+\alpha_{i}\alpha_{k\overline{j}}\big{)}+\frac{2\alpha_{i}\alpha_{k}\alpha_{\overline{j}}}{\alpha^{3}}+\frac{\beta_{ik\overline{j}}}{\beta}-\frac{1}{\beta^{2}}\big{(}\beta_{k}\beta_{i\overline{j}}+\beta_{i}\beta_{k\overline{j}}+\beta_{\overline{j}}\beta_{ik}\big{)}+\frac{2\beta_{i}\beta_{k}\beta_{\overline{j}}}{\beta^{3}}.$
@@
(4.9) $\left\{\begin{array}[]{llll}\tilde{g}_{i\overline{j},k\overline{\ell}}(0)&=&0,\ \ \ \ \mbox{if}\ \{i,k\}\neq\{j,\ell\},\\ \tilde{g}_{i\overline{i},i\overline{i}}(0)&=&-2(\alpha_{i\overline{i}}+\beta_{i\overline{i}})\ =\ -2\,\big{(}1+\delta_{i2}\big{)},\\ \tilde{g}_{i\overline{i},k\overline{k}}(0)&=&\tilde{g}_{i\overline{k},k\overline{i}}(0)\ =\ -(\alpha_{i\overline{i}}\alpha_{k\overline{k}}+\beta_{i\overline{i}}\beta_{k\overline{k}})+|f_{ik}|^{2}\ =\ \left\{\begin{array}[]{ll}-1,\ \mbox{if}\ \{i,k\}=\{1,2\}\ \mbox{or}\ \{2,3\},\\ \ \,1,\ \ \mbox{if}\ \{i,k\}=\{1,3\}.\end{array}\right.\end{array}\right.$
@@
(4.18) $g_{i\overline{i},k\overline{k}}(0)\ =\ \left\{\begin{array}[]{lll}-1+\delta_{i2},\ \ \ \ \ \mbox{if}\ \{i,k\}=\{1,2\}\ \mbox{or}\ \{2,3\},\\ \ \,1-\delta_{i1}\delta_{k3},\ \ \mbox{if}\ \{i,k\}=\{1,3\}.\\ \end{array}\right.$
@@
[…]
@@
(4.28) $\nabla^{b}T=0\ \Longleftrightarrow\ dT^{j}_{ik}=\sum_{r}\big{(}\theta^{b}_{ir}T^{j}_{rk}+\theta^{b}_{kr}T^{j}_{ir}-\theta^{b}_{rj}T^{r}_{ik}\big{)},$
When $g_{i\overline{j}}=\delta_{ij}$ at the origin 00, then by (4.27) the BTP condition at 00 is given by
@@
For (4.29), the right hand side is zero because, for each of these three terms, one of the two factors is zero when $r$ is $2$ or not $2$. Its left hand side at 00 is equal to $(g_{k\overline{j},i\ell}-g_{i\overline{j},k\ell})$, which is zero by the last equality in (4.11). For (4.30), the left hand side at the origin is given by
@@
In the Appendix, we will also verify that the Wallach threefold $(X^{3},g)$ as a Riemannian manifold is Einstein (with constant Ricci curvature $3$) and has non-negative sectional curvature. The metric $g$ is a boundary point of the region of positively curved metrics on $X^{3}$ discovered by Wallach [31].

### v2

@@ 3. The Fano case
(3.1) $\left\{\begin{array}[]{lll}\Theta^{b}_{22}\ =\ dx+\alpha\overline{\alpha},\ \ \ \ \ \Theta^{b}_{33}\ =\ dy-\alpha\overline{\alpha},\\ \Theta^{b}_{23}\ =\ d\alpha-x\alpha-\alpha y,\ \ \ \Theta^{b}_{12}\ =\ \Theta^{b}_{13}\ =\ 0,\\ \Theta^{b}_{11}\ =\ dx+dy\ =\ \Theta^{b}_{22}+\Theta^{b}_{33}.\end{array}\right$.
@@
It follows from [ZZ24+, Theorem 1.1] that the BTP condition $\nabla^{b}T=0$ implies
@@
That is, $\tilde{\omega}$ is a Kähler-Einstein metric with positive Ricci curvature $2$, hence $M^{3}$ is a Fano threefold. Denote by $E$ and $F$ the $C^{\infty}$ complex vector bundles on $M$ whose fibers are $E_{x}={\mathbb{C}}\{e_{2}(x),e_{3}(x)\}$ and $F_{x}={\mathbb{C}}e_{1}(x)$, respectively. They are both globally defined since $E$ is the eigenspace of $B$ corresponding to the eigenvalue 0, and $F$ is the orthogonal complement of $E$ in the holomorphic tangent bundle $T_{\!M}=T^{1,0}M$. For any $i\in\{2,3\}$ and any $j$, we have
@@
(3.7) $0\rightarrow E\rightarrow T_{\!M}\rightarrow L\rightarrow 0.$
@@
Recall that the index of a Fano manifold $X^{n}$ is the largest positive integer $r$ so that $K^{\!-\!1}_{\!X}=A^{\otimes r}$ for an ample line bundle $A$. It is necessarily less than or equal to $n+1$, where $r=n+1$ if and only if $X={\mathbb{P}}^{n}$ and $r=n$ if and only if $X={\mathbb{Q}}^{n}$, the smooth quadratic hypersurface in ${\mathbb{P}}^{n+1}$. Fano manifolds satisfying r=n1r=n-1 are called del Pezzo manifolds, which are classified by Fujita [Fuj90] as one of the following seven types according to their degree $d$, which is the self intersection number $A^{n}$:
@@ 7. $(7)$
For $n=3$, del Pezzo threefolds were classified by Iskovskikh [Isk77] earlier, and in Table 12.2 of [IP99] we can find the third betti number $b_{3}$, hence the Euler number $c_{3}=4-b_{3}$ of del Pezzo threefolds of degree $1\leq d\leq 5$:
@@
If $M^{3}$ is a del Pezzo threefold of degree $d$, then $c_{1}^{3}=8d$, hence the equality (3.10) yields $c_{3}=12-d$. This rules out the possibility of $1\leq d\leq 5$ by (3.9). The case ${\mathbb{P}}^{3}\#\overline{{\mathbb{P}}^{3}}$ of degree $d=7$ has Euler number $c_{3}=6$, which is not equal to 12712-7. Similarly, for the case ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ of degree $d=6$, its Euler number $c_{3}=8$ is not equal to 12612-6, which indicates that neither can be $M^{3}$. Therefore only two possibilities are left, namely, $M^{3}$ is either the flag threefold ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$ or ${\mathbb{P}}^{3}$.
@@
To see this, let $e$ be a local special frame of $(M^{3},g)$, with dual coframe $\varphi$. Let $s$ be a local holomorphic frame of $L$. Since $\xi$ is a $L$-valued 11-form, we have $\xi=\psi\otimes s$ where $\psi$ is a nowhere-zero local holomorphic $1$-form. The kernel of the map given by $\xi$ is $E$, so $\psi=f\varphi_{1}$ for some local smooth function $f$, which is nowhere zero. By the structure equation, it follows that $d\varphi_{1}=-(x+y)\varphi_{1}-\varphi_{2}\varphi_{3}$, so the holomorphicity of $\psi$ gives us
@@
(3.11) $\partial\overline{\partial}\log(|f|^{2}\!\parallel\!\!s\!\!\parallel^{2})=0.$
@@
Assume that $M^{3}={\mathbb{P}}^{3}$. In this case $L={\mathcal{O}}(2)$. It follows from [Sch80] for instance that $\Omega(2)$ admits a trivial line subbundle, which corresponds to a nowhere zero holomorphic section $\xi\in V:=H^{0}({\mathbb{P}}^{3},\Omega(2))$. $\xi$ determines a surjective bundle map $T_{\mathbb{P}^{3}}\rightarrow L$ so that the exact sequence (3.7) holds. Let $\tilde{\omega}$ be the (scaled) Fubini-Study metric of ${\mathbb{P}}^{3}$ with Ricci curvature $2$. It has constant holomorphic sectional curvature $1$. Let $Z_{0}\!:\!Z_{1}\!:\!Z_{2}\!:\!Z_{3}]$ be the standard unitary homogeneous coordinate of ${\mathbb{P}}^{3}$. In the coordinate neighborhood $U_{0}=\{Z_{0}\neq 0\}$, let $z_{i}=\frac{Z_{i}}{Z_{0}}$, $1\leq i\leq 3$ and it follows that
@@
where $c>0$ is a constant. We want to derive at a contradiction. For each $1\leq i\leq 3$, write $\ell_{i}=\ell_{i}^{(0)}+\ell_{i}^{(1)}$ where $\ell_{i}^{(0)},\ \ell_{i}^{(1)}$ are respectively the degree 0 and degree $1$ part of $\ell_{i}$. By looking at the degree $4$ part of the left hand side of (3.12), we get
@@
Now we are left with the only possibility for a compact balanced BTP threefold $(M^{3},g)$ with $B$ rank equal to $1$: the flag threefold $X^{3}={\mathbb{P}}(T_{{\mathbb{P}}^{2}})$. Note that $X^{3}$ is the hypersurface in $N^{4}={\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$ defined by $Z_{0}W_{0}+Z_{1}W_{1}+Z_{2}W_{2}=0$, where $Z$, $W$ are the standard unitary homogeneous coordinate of the two factors of $N$. For $i=1,2$, denote by $\pi_{i}:X^{3}\rightarrow{\mathbb{P}}^{2}$ the restriction on $X$ of the projection map from $N$ onto its $i$-th factor. The Picard group $\mbox{Pic}(X)\cong{\mathbb{Z}}^{\oplus 2}$ is generated by $L_{1}$ and $L_{2}$, where $L_{i}=\pi_{i}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)$, and the anti-canonical line bundle of $X$ is $K_{\!X}=2L=2(L_{1}+L_{2})$. Here for convenience we used the additive notation for line bundles. The Kähler-Einstein metric $\tilde{\omega}$ on $X$ is the restriction of the product of Fubini-Study metric, and we have
@@
Let $X^{3}$ be the flag threefold ${\mathbb{P}}(T_{{\mathbb{P}}^{2}})$, $\Omega$ its holomorphic cotangent bundle, and $L$ be the line bundle $\pi_{1}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)\otimes\pi_{2}^{\ast}{\mathcal{O}}_{{\mathbb{P}}^{2}}(1)$. Then we have
@@
$0\rightarrow T_{X|{\mathbb{P}}^{2}}\rightarrow T_{X}\rightarrow\pi_{1}^{\ast}T_{{\mathbb{P}}^{2}}\rightarrow 0.$
@@
(3.15) $0\rightarrow{\mathcal{O}}_{X}\rightarrow\pi_{1}^{\ast}\Omega_{{\mathbb{P}}^{2}}\!\otimes\!L\rightarrow T_{X|{\mathbb{P}}^{2}}\rightarrow 0.$
@@
where $Z$, $W$ are unitary homogeneous coordinate on the two factors of $N={\mathbb{P}}^{2}\times{\mathbb{P}}^{2}$, viewed as column vectors. It is not hard to see that the norm $\parallel\!\!\sigma\!\!\parallel=\frac{1}{2}$ with respect to the Kähler-Einstein metric $\tilde{\omega}$ of $X$. Consider the Hermitian metric $g$ on $X$ with Kähler form $\omega=\tilde{\omega}-\sigma$, which is clearly a homogeneous Hermitian metric on $X$. We will call the Hermitian manifold $(X,g)$ the Wallach threefold from now on, to honor the influential work [Wal72] in geometry.
We will verify in the next section that $(X,g)$ is indeed balanced and BTP. We will also show that its Chern connection has non-negative bisectional curvature and positive holomorphic sectional curvature, and all three Ricci tensors of the Chern connection are positive. The sectional curvature of the Levi-Civita connection of $g$ is non-negative, and the Levi-Civita connection has constant Ricci curvature $3$, thus $g$ lies in the boundary of the set of metrics with positive sectional curvature discovered by Wallach in [Wal72].
Note that homogeneous metrics on $X$ with positive sectional curvature, which are all Hermitian as observed by Wallach in [Wal72], form a moduli which depends on three real parameters. After scaling, these metrics form a peculiar planer region (see for example Figure 1 in [BM15]). It is not clear which metric in the set is the ‘best’ amongst its peers.
@@ 4. The Wallach threefold
Fix any point $p\in X=SU(3)/T^{2}$. Note that for any $A\in SU(3)$, the map $([Z],[W])\mapsto([AZ],[\overline{A}W])$ is an isometry on $(X,g)$. So without loss of generality, we may assume that p=([1:0:0],[0:0:1])p=([1\!\!:\!0\!:\!0],[0\!:\!0\!:\!\!1]). Consider the neighborhood $U_{02}=\{Z_{0}\neq 0\}\times\{W_{2}\neq 0\}$ in $N$, with local holomorphic coordinate $(z_{1},z_{2},w_{0},w_{1})$ where $z_{i}=\frac{Z_{i}}{Z_{0}}$, $i=1,2$, and $w_{j}=\frac{W_{j}}{W_{2}}$, $j=0,1$. Within $U_{02}$, the hypersurface $X$ is defined by
@@
(4.4) $\tilde{g}_{i\overline{j}}=\frac{\alpha_{i\overline{j}}}{\alpha}-\frac{\alpha_{i}\alpha_{\overline{j}}}{\alpha^{2}}+\frac{\beta_{i\overline{j}}}{\beta}-\frac{\beta_{i}\beta_{\overline{j}}}{\beta^{2}},\ \ \ \ \ \ 1\leq i,j\leq 3,$
@@
(4.5) $\tilde{g}_{i\overline{j},k}=-\frac{1}{\alpha^{2}}\big(\alpha_{k}\alpha_{i\overline{j}}+\alpha_{i}\alpha_{k\overline{j}}\big)+\frac{2\alpha_{i}\alpha_{k}\alpha_{\overline{j}}}{\alpha^{3}}+\frac{\beta_{ik\overline{j}}}{\beta}-\frac{1}{\beta^{2}}\big(\beta_{k}\beta_{i\overline{j}}+\beta_{i}\beta_{k\overline{j}}+\beta_{\overline{j}}\beta_{ik}\big)+\frac{2\beta_{i}\beta_{k}\beta_{\overline{j}}}{\beta^{3}}.\ \ \$
@@
[…]
@@
[…]
@@
$=$ $|X_{2}Y_{1}|^{2}+|X_{2}Y_{3}|^{2}+|X_{1}\overline{Y}_{1}+X_{2}\overline{Y}_{2}|^{2}+|X_{1}\overline{Y}_{1}-X_{3}\overline{Y}_{3}|^{2}+|X_{2}\overline{Y}_{2}+X_{3}\overline{Y}_{3}|^{2}\ \geq\ 0.$
@@
(4.28) $\nabla^{b}T=0\ \Longleftrightarrow\ dT^{j}_{ik}=\sum_{r}\big(\theta^{b}_{ir}T^{j}_{rk}+\theta^{b}_{kr}T^{j}_{ir}-\theta^{b}_{rj}T^{r}_{ik}\big),\$
When $g_{i\overline{j}}=\delta_{ij}$ at the origin 0, then by (4.27) the BTP condition at 0 is given by
@@
For (4.29), the right hand side is zero because, for each of these three terms, one of the two factors is zero when $r$ is $2$ or not $2$. Its left hand side at 0 is equal to $(g_{k\overline{j},i\ell}-g_{i\overline{j},k\ell})$, which is zero by the last equality in (4.11). For (4.30), the left hand side at the origin is given by
@@
In the Appendix, we will also verify that the Wallach threefold $(X^{3},g)$ as a Riemannian manifold is Einstein (with constant Ricci curvature $3$) and has non-negative sectional curvature. The metric $g$ is a boundary point of the region of positively curved metrics on $X^{3}$ discovered by Wallach [Wal72].